# Contrary Inferences for Classical Histories within the Consistent Histories Formulation of Quantum Theory

### v3

@@ Acknowledgements
P.W. thanks Fay Dowker and Rafael Sorkin for early discussions and acknowledges support by the UK Hub in Quantum Computing and Simulation with funding from the EPSRC grant EP/T001062/1. A.Z. acknowledges support by the UNAM Postdoctoral Program (POSDOC).
@@ A.1 The consistent histories formalism
Here again we use the example of a non-relativistic particle for simplicity. The decoherence functional is a complex-valued, bi-linear function defined on the space of all (fine-grained or coarse-grained) histories $h_{i},h_{j}$. The integral is defined over the paths $x^{i}(t)$ in the coarse-grained histories $h_{i}$ and $h_{j}$ respectively. The paths start from $x_{0}$ at $t_{0}$ and end at $x_{f}$, and are defined within the time interval $[t_{0},t_{f}]$, while $\rho(x_{0}^{i},x_{0}^{j})$ is the initial density matrix; the $\delta$-function ensures that histories with different end-point decohere (we can mathematically think of having a final-time position measurement).
@@ Appendix B Quasi-classical and macroscopic approximation
The quantum state (17) satisfies trivially the boundary conditions of the infinite square well since any linear combination of the eigenstates will satisfy them. To take the quasi-classical and macroscopic limit, we have to select the parameters such that $1\ll{\sigma_{0}}\ll{n_{0}}$, see ref. [47] and perform the summations within a range $[n_{0}-n_{f},n_{0}+n_{f}]$ where $n_{f}$ is properly chosen so that the summation converges. In fact, because the energy quantum number $n_{0}$ is large, the summations can also be turned to integrals in the corresponding interval. Moreover, the wave function in the momentum space defined as:
@@
    The classical trajectory of a free particle in 1-dimension, with momentum $p_{0}$, starting position $x_{0}=-L/2$ moving between two walls at ±$\pm L$, is given by
    $x_{cl}(t)=\begin{cases}&x_{\rightarrow}^{(n)}\equiv-2(2n)L+(x_{0}+\frac{p_{0}}{M}t),\quad t_{2n}\leq t\leq t_{2n+1},\ n=0,1,2...\\ &x_{\leftarrow}^{(n)}\equiv 2(2n-1)L-(x_{0}+\frac{p_{0}}{M}t),\quad t_{2n-1}\leq t\leq t_{2n}\ n=1,2,...\end{cases}$ (41)
@@
[…]
@@
[…]
    and is plotted in Figure 7(c). The uncertainty relation for some classical periods is close to /2Planck-constant-over-$\hbar/2$ but after a few periods this picture is lost. The fact that the uncertainty principle does not satisfy the minimal relation (contrary to what happens for the harmonic oscillator) is well known in the case of the coherent states of the infinite square well e.g. [47].
We end this appendix by commenting how we re-scale the variables to make them dimensionless for the numerical calculations presented in this paper. Every position variable ($x$, $\sigma_{0}$, etc.) is defined as a ratio of $L$. The momentum is defined as a ratio of the initial momentum. Finally, the time is defined as a ratio of the classical period $T_{cl}$ corresponding to the well $[-L,L]$.

### v4

@@ Acknowledgements
P.W. thanks Fay Dowker and Rafael Sorkin for early discussions and acknowledges support by the UK Hub in Quantum Computing and Simulation with funding from the EPSRC grant EP/T001062/1. A.Z. acknowledges support by the UNAM Postdoctoral Program (POSDOC) and by the National Science Centre in Poland under the research grant Maestro (2021/42/A/ ST2/00356).
@@ A.1 The consistent histories formalism
Here again we use the example of a non-relativistic particle for simplicity. The decoherence functional is a complex-valued, bi-linear function defined on the space of all (fine-grained or coarse-grained) histories $h_{i},h_{j}$. The integral is defined over the paths $x^{i}(t)$ in the coarse-grained histories $h_{i}$ and $h_{j}$ respectively. The paths start from $x_{0}$ at $t_{0}$ and end at $x_{f}$, and are defined within the time interval $t_{0},t_{f}]$, while $\rho(x_{0}^{i},x_{0}^{j})$ is the initial density matrix; the $\delta$-function ensures that histories with different end-point decohere (we can mathematically think of having a final-time position measurement).
@@ Appendix B Quasi-classical and macroscopic approximation
The quantum state (17) satisfies trivially the boundary conditions of the infinite square well since any linear combination of the eigenstates will satisfy them. To take the quasi-classical and macroscopic limit, we have to select the parameters such that $1\ll{\sigma_{0}}\ll{n_{0}}$, see ref. [47] and perform the summations within a range $n_{0}-n_{f},n_{0}+n_{f}]$ where $n_{f}$ is properly chosen so that the summation converges. In fact, because the energy quantum number $n_{0}$ is large, the summations can also be turned to integrals in the corresponding interval. Moreover, the wave function in the momentum space defined as:
@@
    ${\langle x\rangle}=\int\limits_{-L}^{L}{dx}{{\bar{\Psi}}}(x,t)x\Psi(x,t)$ (40)
    The classical trajectory of a free particle in 1-dimension, with momentum $p_{0}$, starting position $x_{0}=-L/2$ moving between two walls at $\pm L$, is given by
@@
    ${\langle p\rangle}=-i\hbar\int\limits_{-L}^{L}{dx}{\bar{\Psi}}(x,t)\frac{{\partial\Psi(x,t)}}{{\partial x}}$ (42)
@@
    ${\langle{{p^{2}}}\rangle}={\hbar^{2}}\int\limits_{-L}^{L}{dx}{\bar{\Psi}}(x,t)\frac{{{\partial^{2}}\Psi(x,t)}}{{\partial{x^{2}}}}$ (45)
    and is plotted in Figure 7(c). The uncertainty relation for some classical periods is close to $\hbar/2$ but after a few periods this picture is lost. The fact that the uncertainty principle does not satisfy the minimal relation (contrary to what happens for the harmonic oscillator) is well known in the case of the coherent states of the infinite square well e.g. [47].
We end this appendix by commenting how we re-scale the variables to make them dimensionless for the numerical calculations presented in this paper. Every position variable ($x$, $\sigma_{0}$, etc.) is defined as a ratio of $L$. The momentum is defined as a ratio of the initial momentum. Finally, the time is defined as a ratio of the classical period $T_{cl}$ corresponding to the well [L,L][-L,L].